# Primitive-Root Determinant Densities over Prime Fields and Implications for PRIM-LWE

### 11 Mar 2026 v1

@@ Abstract
The prim-lwe problem, introduced by Sehrawat, Yeo, and Desmedt (Theoretical Computer Science, 886 (2021)), is a variant of the Learning with Errors problem in which the secret matrix is required to have a primitive-root determinant. The dimension-uniform reduction constant is $c(p)=\inf_{m\geq 1}c_{m}(p)$, where $c_{n}(p)$ is the exact density of $n\times n$ matrices over $\mathbb{F}_{p}$ with primitive-root determinant. Sehrawat, Yeo, and Desmedt asked whether $\inf_{p\text{ prime}}c(p)=0$, observing that an affirmative answer would follow from the conjectural infinitude of primorial primes.
We resolve this question unconditionally using only Dirichlet’s theorem and Mertens’ product formula, entirely bypassing the primorial-prime hypothesis. We further establish the sharp order
@@
prove that $c(p)$ possesses a continuous purely singular limiting distribution over the primes with support exactly $[0,1/2]$, and derive explicit lower bounds on $c(q)$ for primes of cryptographic interest, parameterized solely by the number $\omega(q-1)$ of distinct prime factors of q1q-1. These bounds apply to every prime $q$ whose predecessor q1q-1 has controlled factorization structure, as measured by the number $\omega(q-1)$ of distinct prime factors; this includes many common NTT-friendly moduli, though NTT-friendliness alone does not imply the needed bound. For the NIST-standardized moduli $q=3329$ and $q=8380417$, the dimension-uniform expected rejection-sampling overhead $1/c(q)$ in the prim-lwe reduction is at most $2.17$ and $3.42$, respectively. As a simple conservative explicit bound, for any prime $q>2^{30}$ the dimension-uniform expected rejection-sampling overhead $1/c(q)$ is at most $1.79\log q$. On the other hand, our results show that the worst-case overhead among primes $p\leq x$ is of order $\Theta(\log\log x)$, and in particular $1/c(q)=O(\log\log q)$ pointwise.
@@ 1 Introduction
Let $p$ be a prime and $M_{n}(\mathbb{F}_{p})$ denote the ring of $n\times n$ matrices over $\mathbb{F}_{p}$. Following [1], a matrix $M\in M_{n}(\mathbb{F}_{p})$ has primitive-root determinant if $\det(M)$ generates $\mathbb{F}_{p}^{*}$. The fraction of such matrices is
$c_{n}(p)\;=\;\frac{\varphi(p-1)}{p-1}\;\prod_{j=1}^{n}\Bigl(1-\frac{1}{p^{j}}\Bigr),$
and since $c_{m}(p)$ is decreasing in $m$, the dimension-uniform constant governing the reduction from decision-LWE to decision-prim-lwe is
$c(p)\;:=\;\inf_{m\geq 1}c_{m}(p)\;=\;\frac{\varphi(p-1)}{p-1}\;\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{p^{j}}\Bigr).$
Sehrawat, Yeo, and Desmedt [1, Remark 3(ii)] observed that $\inf_{p}c(p)=0$ would follow from the conjectural infinitude of primorial primes and asked whether this can be proved unconditionally.
@@
1. Unconditional vanishing (Theorem 3). We prove $\inf_{p\text{ prime}}c(p)=0$ using only Dirichlet’s theorem on primes in arithmetic progressions and Mertens’ third theorem. The key observation is that for each primorial $N_{k}=\prod_{i=1}^{k}p_{i}$, Dirichlet’s theorem supplies infinitely many primes $p\equiv 1\pmod{N_{k}}$, and for any such prime the totient ratio $\varphi(p-1)/(p-1)$ is bounded by the Mertens product $\prod_{i=1}^{k}(1-1/p_{i})\sim e^{-\gamma}/\!\log p_{k}\to 0$. This bypasses the primorial-prime hypothesis entirely.
@@
The matching lower bound follows from two observations: if $p\leq x$ is an odd prime, then $\omega(p-1)\leq m(x):=\max\{k:N_{k}\leq x\}$, and Mertens’ theorem converts this into a $1/\!\log\log x$ lower bound for $\varphi(p-1)/(p-1)$; the additional factor $\prod_{j\geq 1}(1-p^{-j})\geq P_{3}$ is handled by Theorem 8. Hence the decay to zero is completely quantified and extremely slow. To put this rate in perspective, $\log\log(2^{128})\approx 4.5$ and $\log\log(2^{4096})\approx 7.95$, so the asymptotic lower bound $\min_{p\leq x}c(p)\gg 1/\!\log\log x$ already guarantees that the minimum cannot drop faster than the reciprocal of a slowly growing function. Nonetheless, the implied constant is modest (roughly $P_{3}e^{-\gamma}\approx 0.31$), so primes $p$ whose predecessor p1p-1 is divisible by many small primes can still have $c(p)<0.1$ well within practical ranges. The concrete impact on prim-lwe therefore depends on the factorization structure of the chosen modulus; the explicit bounds in Section 6 show that standard cryptographic moduli with controlled factorization incur only a small constant-factor overhead. These implied constants should in principle be capable of being made explicit by combining explicit estimates for the Mertens product and for $\vartheta(x)$ with an explicit least-prime-in-an-arithmetic-progression bound; see Remark 3. We do not pursue that bookkeeping here.
3. Limiting distribution (Theorems 6 and 7, Corollary 2). The quantity $c(p)$ possesses a continuous purely singular limiting distribution over the primes, with support exactly $0,\tfrac{1}{2}]$. If $G$ denotes the distribution function of this law, then
@@
the set of primes with $c(p)\leq\tau$ has relative density $G(\tau)$, and $G(\tau)\to 0$ as $\tau\downarrow 0$.
The existence and continuity of $G$ follow from the shifted-prime Erdős–Wintner framework for additive functions evaluated at p1p-1; for the relevant classical background see Tenenbaum [2] and Deshouillers–Hassani [3]. The full-support conclusion on $0,\tfrac{1}{2}]$, which to our knowledge has not appeared explicitly in the literature, combines the left non-differentiability result of [3, Theorem 1.1] with the density of $\{\varphi(m)/m:m\text{ even}\}$ in $[0,1/2]$ (Lemma 3). Our novel contribution is the transfer of this classical law to $c(p)$ via a transport lemma (Lemma 4), which establishes that the dimension-uniform prim-lwe reduction constant shares the same limiting distribution.
@@
The general $\omega(q-1)$-based bound is immediate from Theorem 8 and gives factors at most $5.36$ and $7.82$ for the NIST-standardized moduli $q=3329$ used in ML-KEM [4] and $q=8380417$ used in ML-DSA [5], respectively. The sharper NIST figures use the exact factorizations recorded in Section 6.
Remark 1.
The constant $P_{3}$ may also be written in standard $q$-series notation as
$P_{3}=\left(\frac{1}{3};\frac{1}{3}\right)_{\infty},\qquad(a;q)_{\infty}:=\prod_{m=0}^{\infty}(1-aq^{m}).$
@@
A related distribution problem was studied by Aivazidis and Sofos [6], who considered the density of maximal-order elements in $\mathrm{GL}_{n}(q)$ as $q$ varies over prime powers and proved corresponding distribution laws. Their setting is different from ours: we work over prime fields and with the primitive-root determinant density arising in prim-lwe. On the distributional side, Deshouillers and Hassani [3, Proposition 2.1 and equation (2.1)] proved limiting-distribution laws for $\varphi(p-1)/(p-1)$ in the residue classes $p\equiv 1\pmod{m}$, including continuity, and [3, Theorem 1.1] established left non-differentiability of the distribution function at every value $\varphi(m)/m$ with $m$ even. The analogous support question for the classical distribution of $\varphi(n)/n$ over the integers remains open; see [3].
@@ 2 Background and the Open Question
exists and is strictly positive for each fixed $p$. The quantity $c_{n}(p)$ is the exact acceptance probability in dimension $n$. Since $c_{m}(p)$ is decreasing in $m$, the dimension-uniform constant that enters the prim-lwe reduction is
$\inf_{m\geq 1}c_{m}(p)=\lim_{m\to\infty}c_{m}(p)=c(p).$
@@
by Mertens’ theorem, which would imply $\inf_{p}c(p)=0$. They left open whether this can be proved unconditionally.
@@ 3 Unconditional Proof that $\inf_{p}c(p)=0$
Remark 2.
The key advantage of this argument over the approach suggested in [1] is that we require only a prime $p\equiv 1\pmod{N_{k}}$—an infinite supply of which is guaranteed by Dirichlet’s theorem—rather than a primorial prime $p=N_{k}+1$, whose infinitude remains a longstanding open problem; see [10] for computational results on primorial primes.
@@ 5 The Limiting Distribution of $c(p)$
We now turn to the distributional theory of $c(p)$ over the primes. A related density problem in $\mathrm{GL}_{n}(q)$, with $q$ ranging over prime powers, was studied by Aivazidis and Sofos [6], who proved a distribution law for the density of maximal-order elements. Our setting is different: we work over prime fields and with the prim-lwe density $c(p)$. Recall that
$f(n):=\log\frac{n}{\varphi(n)}=\sum_{\ell\mid n}\log\frac{\ell}{\ell-1},\qquad\frac{\varphi(p-1)}{p-1}=e^{-f(p-1)},$
and hence the problem sits inside the shifted-prime Erdős–Wintner framework [13] for additive functions on $p+a$ with fixed $a\neq 0$—for this viewpoint, see Tenenbaum [2]. Its singularity is also essentially classical, ultimately resting on the Jessen–Wintner theory of infinite convolutions [14]: in the integer setting, pure singularity under the condition
$\sum_{\ell\ \text{prime}}\frac{f(\ell)^{2}}{\ell}<\infty$
is due to Erdős [15]; see also [16, Exercise 256]. Tenenbaum [2, p. 258] observes that the same conclusion carries over to the shifted-prime law by adapting the same argument to the shifted-prime characteristic function. The full-support conclusion on $0,\tfrac{1}{2}]$ then follows from Deshouillers and Hassani [3, Theorem 1.1] together with Lemma 3. Our novel contribution in this section is the transport of this classical law to $c(p)$ via Lemma 4, which establishes that the dimension-uniform prim-lwe reduction constant shares the same limiting distribution.
@@ 5.1 The distribution of $\varphi(p-1)/(p-1)$
Historically, the distribution of additive functions on shifted primes $p+a$, with fixed $a\neq 0$, is governed by the shifted-prime analogue of the Erdős–Wintner theorem; see [17, 18, 2] for a concise overview and references. For the specific function needed here, however, one can bypass that general machinery and use the following direct p1p-1-specific input from Deshouillers and Hassani.
Proposition 2 (Deshouillers–Hassani; special case of [3, Proposition 2.1 and equation (2.1)]).
@@
Apply Deshouillers and Hassani [3, Proposition 2.1] to
@@
For a prime $\ell$,
$\log g(\ell)=\log\!\left(1-\frac{1}{\ell}\right)=-\frac{1}{\ell}+O\!\left(\frac{1}{\ell^{2}}\right),$
so, writing $h(\ell):=\log g(\ell)$ for each prime $\ell$, we have $|h(\ell)|=|\log(1-1/\ell)|\leq\log 2<1$ for every prime $\ell$. Therefore the third Erdős–Wintner series $\sum_{|h(\ell)|>1}1/\ell$ is vacuous, while
$\frac{h(\ell)}{\ell}=-\frac{1}{\ell^{2}}+O\!\left(\frac{1}{\ell^{3}}\right),\qquad\frac{h(\ell)^{2}}{\ell}=\frac{1}{\ell^{3}}+O\!\left(\frac{1}{\ell^{4}}\right).$
Hence
$\sum_{\ell}\frac{h(\ell)}{\ell}\quad\text{converges absolutely, and}\quad\sum_{\ell}\frac{h(\ell)^{2}}{\ell}\quad\text{converges}.$
Thus the three convergence hypotheses of [3, Proposition 2.1] are satisfied, and there exists a distribution function $G_{m}$ such that the displayed limit holds at every continuity point of $G_{m}$.
@@
Apply Proposition 2 with $m=2$. Since
@@
By Deshouillers and Hassani [3, Theorem 1.1], the distribution function $G$ is not left-differentiable at $x_{m}$. If $x_{m}\notin\operatorname{supp}(\mu_{\varphi})$, then some open interval $I$ containing $x_{m}$ has $\mu_{\varphi}(I)=0$. Hence $G$ is locally constant on $I$, and in particular left-differentiable at $x_{m}$, a contradiction. Therefore $x_{m}\in\operatorname{supp}(\mu_{\varphi})$ for every even $m$.
@@
Theorem 6 completely determines the shifted-prime support: it is exactly $0,\tfrac{1}{2}]$. This does not settle the classical support question for the distribution of $\varphi(n)/n$ over the integers, which remains open; see [3]. The residue-class laws in [3, Proposition 2.1 and equation (2.1)] are consistent with this picture, since for primes $p\equiv 1\pmod{m}$ the corresponding limiting measure is supported in $0,\varphi(m)/m]$.

### 18 Mar 2026 v2

@@ Abstract
The prim-lwe problem, introduced by Sehrawat, Yeo, and Desmedt (Theoretical Computer Science, 886 (2021)), is a variant of the Learning with Errors problem in which the secret matrix is required to have a primitive-root determinant. The dimension-uniform reduction constant is $c(p)=\inf_{n\geq 1}c_{n}(p)$, where $c_{n}(p)$ is the exact density of $n\times n$ matrices over $\mathbb{F}_{p}$ with primitive-root determinant. Sehrawat, Yeo, and Desmedt asked whether $\inf_{p\text{ prime}}c(p)=0$, observing that an affirmative answer would follow from the conjectural infinitude of primorial primes. We resolve this question unconditionally using only Dirichlet’s theorem and Mertens’ product formula, entirely bypassing the primorial-prime hypothesis. We further establish the sharp order
@@
and show that the limiting distribution of $c(p)$ over the primes has support exactly $[0,1/2]$. We have not found this full-support statement in the literature. The law coincides with the classical shifted-prime distribution of $\varphi(p-1)/(p-1)$ via a transport lemma and is moreover continuous and purely singular. We also derive explicit lower bounds on $c(q)$ for primes of cryptographic interest, parameterized solely by the number of distinct prime factors of q1q-1. As a simple conservative explicit bound, for any prime $q>2^{30}$ the expected overhead $1/c(q)$ is at most $1.79\log q$. On the other hand, our results show that the worst-case overhead among primes $p\leq x$ is of order $\Theta(\log\log x)$, and in particular $1/c(q)=O(\log\log q)$ pointwise.
@@ 1 Introduction
Let $p$ be a prime and $M_{n}(\mathbb{F}_{p})$ denote the ring of $n\times n$ matrices over $\mathbb{F}_{p}$. Following [1], a matrix $M\in M_{n}(\mathbb{F}_{p})$ has primitive-root determinant if $\det(M)$ generates $\mathbb{F}_{p}^{*}$. The fraction of such matrices is denoted $c_{n}(p)$; since it is decreasing in $n$, the dimension-uniform constant governing the reduction from decision-LWE to decision-prim-lwe is $c(p):=\inf_{n\geq 1}c_{n}(p)=\lim_{n\to\infty}c_{n}(p)$. The precise formulas are recorded in Section 2. Sehrawat, Yeo, and Desmedt [1, Remark 3(ii)] asked whether $\inf_{p}c(p)=0$ can be proved unconditionally; see Section 2 for the precise statement.
@@
1. Unconditional vanishing (Theorem 3). We prove $\inf_{p\text{ prime}}c(p)=0$ using only Dirichlet’s theorem on primes in arithmetic progressions and Mertens’ third theorem, without assuming the primorial-prime hypothesis; see Section 3.
@@
The matching lower bound follows from two observations: if $p\leq x$ is an odd prime, then $\omega(p-1)\leq m(x):=\max\{k:N_{k}\leq x\}$, and Mertens’ theorem converts this into a $1/\!\log\log x$ lower bound for $\varphi(p-1)/(p-1)$; the additional factor $\prod_{j\geq 1}(1-p^{-j})\geq P_{3}$ is handled by Theorem 8. Hence the decay to zero is completely quantified and extremely slow. The implied constant is $\approx 0.31$, so primes $p$ whose predecessor p1p-1 is divisible by many small primes can still have $c(p)<0.1$, well within practical ranges. The explicit bounds in Section 6 show that standard cryptographic moduli with controlled factorization incur only a small constant-factor overhead; see Remark 1 for effectivity.
3. Limiting distribution (Theorems 6 and 7, Corollary 2). The quantity $c(p)$ possesses a continuous limiting distribution over the primes whose support is exactly $0,\tfrac{1}{2}]$; the law is moreover purely singular. If $G$ denotes the distribution function, then
@@
the set of primes with $c(p)\leq\tau$ has relative density $G(\tau)$, and $G(\tau)\to 0$ as $\tau\downarrow 0$. The existence and continuity of $G$ follow from the shifted-prime Erdős–Wintner framework for additive functions evaluated at p1p-1 [2]; for the relevant classical background see Tenenbaum [3] and Deshouillers–Hassani [4]. Pure singularity is classical in the integer case (Erdős [5]) and carries over to shifted primes by the same method, as noted by Tenenbaum [3, p. 263]; we include a self-contained proof for the reader’s convenience. Our new contributions on the distributional side are twofold. First, we establish that the support of the limiting law is exactly $0,\tfrac{1}{2}]$ by combining the left non-differentiability result of [4, Theorem 1.1] with the density of $\{\varphi(m)/m:m\text{ even}\}$ in $[0,1/2]$ (Lemma 4). Second, we transfer the classical shifted-prime law for $\varphi(p-1)/(p-1)$ to the prim-lwe density $c(p)$ via a transport lemma (Lemma 5).
@@
The general $\omega(q-1)$-based bound is immediate from Theorem 8; the sharper NIST figures use the exact factorizations recorded in Section 6.
@@
A related distribution problem was studied by Aivazidis and Sofos [6], who considered the density of maximal-order elements in $\mathrm{GL}_{n}(q)$ as $q$ varies over prime powers and proved corresponding distribution laws. Their setting is different from ours: we work over prime fields and with the primitive-root determinant density arising in prim-lwe. On the distributional side, Deshouillers and Hassani [4, Proposition 2.1 and equation (2.1)] proved limiting-distribution laws for $\varphi(p-1)/(p-1)$ in the residue classes $p\equiv 1\pmod{m}$, including continuity, and [4, Theorem 1.1] established left non-differentiability of the distribution function at every value $\varphi(m)/m$ with $m$ even. The analogous support question for the classical distribution of $\varphi(n)/n$ over the integers remains open; see [4].
@@ 2 Background and the Open Question
exists and is strictly positive for each fixed $p$. The quantity $c_{n}(p)$ is the exact acceptance probability in dimension $n$. Since $c_{n}(p)$ is decreasing in $n$, the dimension-uniform constant that enters the prim-lwe reduction is
$\inf_{n\geq 1}c_{n}(p)=\lim_{n\to\infty}c_{n}(p)=c(p).$
@@
by Mertens’ theorem, which would imply $\inf_{p}c(p)=0$. They left open whether this can be proved unconditionally; see [8] for computational results on primorial primes.
@@ 5 The Limiting Distribution of $c(p)$
We now turn to the distributional theory of $c(p)$ over the primes. The classical existence and continuity of the limiting distribution of $\varphi(p-1)/(p-1)$ over primes follow from the shifted-prime Erdős–Wintner framework; for the shift a=1a=-1 used here, direct references are Tenenbaum [3, pp. 257–258] and Deshouillers–Hassani [4].
Our contributions in this section are:
  1. (i)
    The full-support conclusion: the limiting law has support exactly $0,\tfrac{1}{2}]$ (Theorem 6), combining the left non-differentiability result of [4, Theorem 1.1] with the density of $\{\varphi(m)/m:m\text{ even}\}$ in $[0,1/2]$ (Lemma 4).
  2. (ii)
    The transport of the classical shifted-prime law to the prim-lwe density $c(p)$ via Lemma 5, establishing that $c(p)$ shares the same limiting distribution as $\varphi(p-1)/(p-1)$.
We supply a self-contained proof of the shifted-prime pure singularity in Lemma 9.
@@ 5.1 The distribution of $\varphi(p-1)/(p-1)$
One can bypass the general Erdős–Wintner machinery and use the following direct p1p-1-specific input from Deshouillers and Hassani.
Lemma 3 (Erdős–Wintner hypotheses for the totient function).
Let $h$ be the strongly additive function $h(n)=\log(n/\varphi(n))=-\log(\varphi(n)/n)$. For every prime $\ell$,
$0<h(\ell)=\log\frac{\ell}{\ell-1}\leq\log 2<1,\qquad h(\ell)=\frac{1}{\ell}+O\!\left(\frac{1}{\ell^{2}}\right).$
Consequently the third Erdős–Wintner series $\sum_{|h(\ell)|>1}1/\ell$ is vacuous, while $\sum_{\ell}h(\ell)/\ell$ and $\sum_{\ell}h(\ell)^{2}/\ell$ both converge. Since $h(\ell)\neq 0$ for every prime $\ell$ and $\sum_{\ell}1/\ell=\infty$, the continuity criterion of Lemma 6 is satisfied.
The same conclusions hold for $g(n)=\varphi(n)/n$ via $\log g(\ell)=-h(\ell)$.
Proposition 2 (Deshouillers–Hassani; special case of [4, Proposition 2.1 and equation (2.1)]).
@@
Apply Deshouillers and Hassani [4, Proposition 2.1] to
@@
By Lemma 3 applied to $\log g(\ell)=-\log(\ell/(\ell-1))$, the three convergence hypotheses of [4, Proposition 2.1] are satisfied, and there exists a distribution function $G_{m}$ such that the displayed limit holds at every continuity point of $G_{m}$.
@@
Apply Proposition 2 with $m=2$. Since every odd prime satisfies $p\equiv 1\pmod{2}$, we have
@@
so the distribution function $G_{2}$ furnished by that proposition is the unconditional limiting distribution of $\varphi(p-1)/(p-1)$ over all primes. In particular, $G_{2}$ coincides with the distribution function denoted $F$ in [4, Theorem 1.1]. Writing $G:=G_{2}$, there exists a continuous distribution function $G\colon\mathbb{R}\to[0,1]$ such that for every $\alpha\in\mathbb{R}$,
@@
By Deshouillers and Hassani [4, Theorem 1.1], the distribution function $G$ is not left-differentiable at $x_{m}$. Deshouillers and Hassani denote this distribution function by $F$; we write it as $G$. If $x_{m}\notin\operatorname{supp}(\mu_{\varphi})$, then some open interval $I$ containing $x_{m}$ has $\mu_{\varphi}(I)=0$. Hence $G$ is locally constant on $I$, and in particular left-differentiable at $x_{m}$, a contradiction. Therefore $x_{m}\in\operatorname{supp}(\mu_{\varphi})$ for every even $m$.
@@
Theorem 6 completely determines the shifted-prime support: it is exactly $0,\tfrac{1}{2}]$. This does not settle the classical support question for the distribution of $\varphi(n)/n$ over the integers, which remains open; see [4]. The residue-class laws in [4, Proposition 2.1 and equation (2.1)] are consistent with this picture, since for primes $p\equiv 1\pmod{m}$ the corresponding limiting measure is supported in $0,\varphi(m)/m]$.
@@ 5.2 From $\varphi(p-1)/(p-1)$ to $c(p)$
The following transport lemma is the mechanism for contribution (ii) listed at the beginning of this section.